OpenStudy (anonymous):

k^2 - 36/k^2 - 16 k + 60 ÷ k^2 - 12 k + 36/k^2 - 6k ⋅ 6k - 60/k^2 + 6k = ?

OpenStudy (anonymous):

@Conqueror @SolomonZelman

OpenStudy (solomonzelman):

I am not going to guess what the denominators and numerators are 500 times. Can you please re-write the question precisely?

OpenStudy (anonymous):

\[\frac{ k^2 - 36 }{ k^2 - 16 k + 60 } \div \frac{ k^2 - 12 k + 6 }{ k^2 - 6 k } \times \frac{ 6 k - 60 }{ k^2 + 6 k } = ?\]

OpenStudy (anonymous):

factorise numerators and denominators wherever possible, rest you can do i suppose, @greenhawk2000

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I need help

OpenStudy (mathmate):

Cannot factorize? Hints: \[a(a+b)=a^2+ab\] \[a^2-b^2=(a+b)(a-b)\] \[(a+b)^2=a^2+2ab+b^2\] \[(a-b)^2=a^2-2ab+b^2\]